Controls Meeting 17-Oct-2013
- Minutes from previous meeting 22-Aug-2013
- FCAL dark room update - John
- Network installation - Elliott, Mark
- Controls installation - Dave B
- new network, moving computers, etc.
- JInventory update - Serguei P
- EPICS update - Hovanes, Nerses, Vanik
- Solenoid improvements - Dave, Mark, Scot, Elliott
- UPS power - Elliott
- Mantis task tracker
- Midas elog
- Future topics
- Experiment Control Supervisor - Vardan, Elliott
9:30AM Thurs 17-Oct-2013 CC A110
9:30am Thurs 31-Oct-2013
Present: Elliott W, Dave B, Hovanes E, Mark S.
See link above for detailed networking plan. We decided to reserve the top left 9 ports for DAQ, the remainder for controls (max ROCs in one area is 9).
- About to test CompactLogix to EPICS communication, ORNL said their driver works.
- Solenoid stuff moved to controls (26) subnet, license servers working, etc.
- PLC database still resides on gluon01 on 36 subnet, now accessible from 26 subnet.
- Dave still needs to update some PLC config information.
- Not clear how many computers needed for PLC support in counting house and hall, Dave will figure this out.
- FCAL PLC installed, being tested.
- FDC or BCAL probably next.
- Need to enter power (Giles), remaining controls equipment, some additional DAQ boards.
- Full role in online not clear, will we get the translation table from the JInventory database?
- Still need to add maintenance functionality for Fernando (we're waiting for his requirements).
Not clear if we need ramp profiles now that basic slew-rate setting is working. It seems that a firmware update caused the problem, went away with another update.
Still need UPS requirements for hall equipment.